# Rotor-Failure-Aware Quadrotors Flight in Unknown Environments

### v1

@@ VI SIMULATION and Real-World Experiments
We evaluate the effectiveness of the proposed method through a combination of simulations and real-world experiments. This section focuses on addressing three critical questions: i) Can the control system handle motor and propeller failures, achieving flight stability without prior knowledge of the failure? (Refer to Sections VI-A and VI-B.) ii) Can the navigation module generate dynamically feasible and collision-free trajectories in cluttered environments, even with unknown obstacles and low-light conditions? (Refer to Section VI-C.) iii) How does the rotor-failure-aware framework perform in the wild? (Refer to Section VI-D.)
The physical parameters of the quadrotor are listed in Table II. For trajectory planning, the unconstrained optimization problem is solved using LBFGS [33]. The NMPC problem is handled by the ACADO [34] toolkit in conjunction with qpOASES [35], operating at a frequency of 200 Hz. The horizon length $N$ and the time step $dt$ are set to 20 and 50 $ms$, respectively. The parameters of the FTC system are specified as $\bm{Q}_{p}=\mathrm{diag}(100,100,600)$, $\bm{Q}_{v}=\mathrm{diag}(5,5,5)$, $\bm{Q}_{q}=\mathrm{diag}(60,60,60,60)$, $\bm{Q}_{w}=\mathrm{diag}(5,5,5)$, $\bm{Q}_{t}=\mathrm{diag}(1,1,1,1)$, $\bm{R}_{N}=\mathrm{diag}(1,1,1,1)$, $\gamma_{M}=0.2$, $\gamma_{P}$ = $0.8$, $\gamma_{Q,0}=0.005\ m/s^{2}$ , $\gamma_{Q,1}=0.005\ m/s^{2}$ , $\gamma_{Q,2}=0.2\ rad/s^{2}$, $\gamma_{Q,3}=0.2\ rad/s^{2}$. The weighting vector $\bm{\lambda}^{p}=[10.0,1.0,10000.0,10000.0]^{T}$ balances the trade-offs among optimization objectives and the safe distance is 0.3 $m$. To quantitatively evaluate the tracking performance, let $\bm{\eta}_{t}=[x_{t},y_{t},z_{t}]^{\top}$ denote the UAV position at time $t$, and $\bm{\eta}_{r}=[x_{r},y_{r},z_{r}]^{\top}$ the corresponding reference position. The root mean square error (RMSE) of trajectory tracking is given by
@@ VI-A2 FTC during Takeoff
Test #1
Hovering
Propeller failure
FDD time ($s$) 0.216 $\bm{0.105}$
@@
Test #2
Hovering
Motor failure
FDD time ($s$) 0.207 $\bm{0.032}$
@@
Test #1
Trajectory tracking
Propeller failure
FDD time ($s$) 0.253 $\bm{0.184}$
@@
Test #2
Trajectory tracking
Motor failure
FDD time ($s$) 0.225 $\bm{0.028}$
@@
Test #3
Takeoff
FDD time ($s$) / 0.195 $\bm{0.020}$
@@
Test #4
Takeoff
FDD time ($s$) / 0.190 $\bm{0.020}$
@@ VI-C Autonomous Indoor Navigation
To fully evaluate the autonomous navigation capability of the proposed planning method, we conduct waypoint navigation experiments in indoor environments. The quadrotor, with one propeller unloaded, undergoes two types of tests: 1) The quadrotor navigates through 6 predefined waypoints while encountering manual-moving obstacles to assess its obstacle avoidance performance in complex environments (see Fig. 11). 2) The quadrotor’s autonomous navigation capability is further tested under low light and messy obstacles to evaluate its performance in a dark environment (see Fig. 12). The total flight time is 66.7 $s$ in Fig. 11, during which the quadrotor operates autonomously with maximum velocity 1.0 $m/s$. With no prior knowledge of the environment apart from the 6 predefined waypoints, the planner efficiently calculates smooth trajectories, enabling the quadrotor autonomous navigation despite an operator moving the obstacles (see the marked obstacles in Fig. 11). The RMSE in flight is 0.08 $m$ and the maximum rotation speed is 10.7 $rad/s$. The computation times of the navigation modules running on the NVIDIA Jetson Orin NX are also analyzed. The average processing time for the LiDAR inertial odometry and the trajectory planner is less than 20 $ms$.
@@ VI-D Autonomous Flight in the Wild
Figure 12: Autonomous waypoint navigation in the low-light environment against propeller-unloading failure.
@@ VII Conclusion
This study highlights the importance of ensuring autonomous navigation and flight stability for quadrotors subject to rotor failures in unknown environments. To address these challenges, we design a composite FDD-based FTC system that promptly acquires information on both motor and propeller failures while mitigating vibrations caused by rotor imbalance, from takeoff through trajectory tracking. In addition, we propose a rotor-failure-aware planning algorithm that formulates an optimization problem incorporating quadrotor dynamic feasibility and obstacle avoidance. Together, these components systematically tackle the problem of achieving collision-free flight in complex environments under rotor failures. Comprehensive simulations and experimental validations confirm the practicality and effectiveness of the proposed method. Despite these contributions, the current system has limitations. In particular, scenarios involving dual rotor loss present severe challenges to controllability that are not addressed in this work. Moreover, strong wind conditions may introduce unpredictable disturbances that degrade control performance. Addressing these scenarios will form an important direction for future research.

### v2

@@ VI SIMULATION and Real-World Experiments
We evaluate the effectiveness of the proposed method through a combination of simulations and real-world experiments. This section focuses on addressing three critical questions: i) Can the control system handle motor and propeller failures, achieving flight stability without prior knowledge of the failure? (Refer to Sections VI-A and VI-B.) ii) Can the navigation module generate dynamically feasible and collision-free trajectories in cluttered environments, even with unknown obstacles and low-light conditions? (Refer to Section VI-C.) iii) How does the rotor-failure-aware framework perform in the wild? (Refer to Section VI-D.) The physical parameters of the quadrotor are listed in Table II. For trajectory planning, the unconstrained optimization problem is solved using LBFGS [35]. The NMPC problem is handled by the ACADO [36] toolkit in conjunction with qpOASES [37], operating at a frequency of 200 Hz. The horizon length $N$ and the time step $dt$ are set to 20 and 50 $ms$, respectively. The parameters of the FTC system are specified as $\bm{Q}_{p}=\mathrm{diag}(100,100,600)$, $\bm{Q}_{v}=\mathrm{diag}(5,5,5)$, $\bm{Q}_{q}=\mathrm{diag}(60,60,60,60)$, $\bm{Q}_{w}=\mathrm{diag}(1,1,1)$, $\bm{Q}_{t}=\mathrm{diag}(1,1,1,1)$, $\bm{R}_{N}=\mathrm{diag}(1,1,1,1)$, $\gamma_{M}=0.2$, $\gamma_{P}$ = $0.8$. The weighting vector $\bm{\lambda}^{p}=[10.0,1.0,10000.0,10000.0]^{T}$ balances the trade-offs among optimization objectives and the safe distance is 0.3 $m$. To quantitatively evaluate the tracking performance, let $\bm{\eta}_{t}=[x_{t},y_{t},z_{t}]^{T}$ denote the UAV position at time $t$, and $\bm{\eta}_{r}=[x_{r},y_{r},z_{r}]^{T}$ the corresponding reference position. The root mean square error (RMSE) of trajectory tracking is given by
@@ VI-A2 FTC during Takeoff
Test #1 Hovering Propeller failure FDD time ($s$) 0.216 $\bm{0.105}$
@@
Test #2 Hovering Motor failure FDD time ($s$) 0.207 $\bm{0.032}$
@@
Test #1 Trajectory tracking Propeller failure FDD time ($s$) 0.253 $\bm{0.184}$
@@
Test #2 Trajectory tracking Motor failure FDD time ($s$) 0.225 $\bm{0.028}$
@@
Test #3 Takeoff FDD time ($s$) / 0.195 $\bm{0.020}$
@@
Test #4 Takeoff FDD time ($s$) / 0.190 $\bm{0.020}$
@@ VI-C Autonomous Indoor Navigation
To fully evaluate the autonomous navigation capability of the proposed planning method, we conduct waypoint navigation experiments in indoor environments. The quadrotor, with one propeller unloaded, undergoes two types of tests: 1) The quadrotor navigates through 6 predefined waypoints while encountering manual-moving obstacles to assess its obstacle avoidance performance in complex environments (see Fig. 11). 2) The quadrotor’s autonomous navigation capability is further tested under low-light conditions and cluttered obstacles to evaluate its performance in a dark environment. [^1] https://sojustfish.github.io/Rotor-Failure-Aware/ The total flight time is 66.7 $s$ in Fig. 11, during which the quadrotor operates autonomously with maximum velocity 1.0 $m/s$. With no prior knowledge of the environment apart from the 6 predefined waypoints, the planner efficiently calculates smooth trajectories, enabling the quadrotor autonomous navigation despite an operator moving the obstacles (see the marked obstacles in Fig. 11). The RMSE in flight is 0.08 $m$ and the maximum rotation speed is 10.7 $rad/s$. The computation times of the navigation modules running on the NVIDIA Jetson Orin NX are also analyzed. The average processing time for the LiDAR inertial odometry and the trajectory planner is less than 20 $ms$.
@@ VII Conclusion
In this study, we underscore the importance of ensuring autonomous navigation and flight stability for a quadrotor experiencing rotor failures in unknown environments. To meet these critical requirements, we design a composite FDD-based FTC system that actively and promptly acquires information on both motor and propeller failures, while mitigating vibrations caused by rotor imbalance from takeoff to trajectory tracking. Furthermore, we propose a rotor-failure-aware planning algorithm that formulates an optimization problem incorporating quadrotor dynamic feasibility and obstacle avoidance. Our approach systematically addresses the challenges of achieving collision-free flight in challenging environments for quadrotors with rotor failures. Comprehensive simulations and experimental validations demonstrate the practicality and effectiveness of the proposed method. Nevertheless, the current system has limitations. In particular, scenarios involving dual rotor loss present severe challenges to controllability that are not yet addressed in this work. Furthermore, strong wind conditions can introduce unpredictable disturbances that degrade control performance. Addressing these scenarios will be an important direction for future research.